What are the number of integer solutions of the equation 7x + 3y = 123, for x, y > 0?

Grade:12th Pass

1 Answers

Chetan Mandayam Nayakar
13 years ago

y=(123-7x)/3 =41-(7x/3), now 7 and 3 are relatively prime. therefore x has to be a multiple of 3 such that y is positive.

from trial and error (x,y) =(3,34),(6,27),(9,20),(12,13),(15,6). thus, there are 5 solutions